IPP Progress Report for the week 2011.07.25 - 2011.07.29

(Up to IPP Progress Reports)

Eugene Magnier

We have been fighting with hardware problems this past week. One machine, ipp030, had a motherboard failure. Cindy replaced it with the motherboard from one of the compute nodes (ippc10). We had many repeated crashes of ippdb03 (backup database server), and have arranged to move its services to another compute node (ippc11). We had a problem with the raid on ippdb00 (nebulous server), compounded by inconsistencies in the reported drive IDs. We had multiple crashes of ipp026, and have removed it from the nebulous list for now. It is unclear to me if this is just a string of bad luck, or if there is an external reason for multiple hardware problems. Cindy et al checked on the load on the PDUs, but there does not seem to be any thing unusual. A/C seems to be fine.

I have been making progress on improving the non-linear fitting for extended sources. I improved the model guess by using a large set of fake data to generate a better set of relationships between observed moments and model parameters. I also improved the speed of the non-linear fitting by (a) using fewer pixels for the fit than for the model subtraction (these had previously been coupled) and (b) allowing for a 1D PSF convolution (saves about a factor of 2).

Chris & I investigated some odd detections reported by MOPS. It seems that these are coming from reflections off vertical structures in the camera close to the focal plane. One set are due to the calcite cells and the second set are (we think) due to the frame around the edge of the focal plane. The resulting artifacts are extremely PSF-like and land within 60 pixels of the edge of the camera (in the second case) or around the edge of the calcite cell. This is one of those annoying cases where we can add a suspect mask bit to warn about the danger, but we probably don't want to exclude all possible detections in this region. Fortunately, these are fairly rare, and only were noticed by MOPS in observations near the Galactic plane. Also fortunately, they should not appear static with respect to the stars because small shifts between the TTI images are doubled for the reflections.

Serge Chastel

  • IPP czar 2 days
  • MOPS czar all week
  • Fixes (IPP failures, database crashes)
  • SAS2 for MOPS

Heather Flewelling

  • liason
    • transferred files to TD
    • helped DF install ipp
  • plots of streak fractions of ThreePi?.
  • investigated cosmic rays for mops
  • addstar
    • fixed it so that 'my' tables get generated when creating a new gpc1-like db.
    • restarted ThreePi? merging
    • send bill instructions for stop/start/restarting addstar
  • czar - unstuck summitcopy/registration several times
  • out sick 2 days
  • vacation day friday

Roy Henderson

  • Only worked for half the week, vacation the rest
  • various discussions about the PSPS changes necessary to host both previous and current versions of 3PI data
  • worked with Mark on the test-loop stuff:
    • implemented some more 'test-mode' code to hard-code certain values if missing from test data
    • problem with renamed extension ('chip' rather than eg XY33)
    • issue with DVO database. Investigated and found missing ID columns in cpm file, eg EXT_ID
  • started work on dvoToMySQL code to investigate duplicates in DVO cpm files discovered before Boston trip:
    • a lot of changes to dvoToMySQL program due to recent changes in other classes
    • code now at stage where I can produce a MySQL Db. Will try it this week
  • added sorely needed comments to Jython ippToPsps code

Mark Huber

  • MD10.V3 work continued - finished reprocessing of exposures through warp, remaining nightly stack reprocessing started and datastore distribution setup. Testing of ppSub automatic choice for convolution direction continues.
  • simtest IPP to PSPS test sample of camera SMFs and generated DVO catalog more complete now.

Bill Sweeney

  • spent most of the week catching up after a 3 week vacation.
  • 2 days as processing czar. Spent endleuss hours repairing faults.
  • talked with Zeliko Ivezic about getting the LSST project some sample PS1 data. Regenerated some sample data for him using the postage stamp server.
  • Kick started the STS processing which got stuck during my absence.
  • started working on new data base tables to support the new video photometry stage.

Chris Waters

  • LAP: rewrote code to deprecate quickstacks. Preliminary test on faked data seemed to work correctly, but a full test with real data still needs to be done.
  • ROC: ran validation code to determine what can be rsynced to the ATRC. Cleaned rsynced data, and now need to correct the files that did not validate correctly.
  • DQstats: multiple bugfixes to correct missing exposures and repeated publishing of data quality statistics to OTIS.
  • SAS: finished SAS processing. Will clean it up once MOPS has received all stamps.
  • MOPS Ghosts: Helped track down positions of ghost detections, and identified that they likely belong to two classes of likely reflections. Wrote script to help identify future ghost objects.